Use the elimination method to solve the following system of equations.
2x + 4y – 3z = –7
3x + y + 4z = –12
x + 3y + 4z = 4

Answer by Alan3354(69443)

Use the elimination method to solve the following system of equations.
2x + 4y – 3z = –7
3x + y + 4z = –12
x + 3y + 4z = 4
====================
This is covered well on google, youtube and wikipedia.